How do I solve this equation 5x^2+3x-2=0 I need to see the steps as I am struggling to understand how to solve this using the quadratic formula
Answer by mananth(16949)

a= 5 , b= 3 , c= -2

b^2-4ac=9+40

b^2-4ac=49
sqrt(49)=7
x=(-b+-sqrt(b^2-4*a*c))/(2*a)

x1=(-b+sqrt(b^2-4*a*c))/(2*a)
x1=(-3+ 7)/10
x1= 0.40

x2=(-3-7) /10
x2= -1.00